OpenStudy (anonymous):

Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form. HELP NEEDED

OpenStudy (asadkarim7):

the parallel lines have same slope/gradient so the gradient would be -1 y=mx+c 2=-1(-4)+c c=-2 y=-x-2